OpenStudy (anonymous):

7+2x=4x-3 solve for X

OpenStudy (jessiegonzales):

x=5

OpenStudy (jessiegonzales):

7+2(5)=4(5)-3 7+10=20-3 17=17 x=5 :)

OpenStudy (anonymous):

try to combine like term

OpenStudy (anonymous):

put x in one side and variable in 1 side